Find the integral curves of the equations dx/yz = dy/zx = dz/xy
1
Expert's answer
2020-07-28T18:07:58-0400

"\dfrac{dx}{yz}=\dfrac{dy}{zx}\newline
\dfrac{dx}{y}=\dfrac{dy}{x}\newline
xdx=ydy\newline
x^2=y^2 + C_1\newline
In\, the\, same\, way:\newline
z^2=y^2+C_2"